OpenStudy (anonymous):

Use the commutative property to rewrite the expression: 5x+(-3y) A.) 3y+(-5x) B.)-3y+5x C.)2xy D.)-15xy

OpenStudy (mathstudent55):

The commutative property of addition states that you can change the order of the numbers in an addition without affecting the result. Thus is usually stated mathematically as: a + b = b + a An example is 5 + 3 = 3 + 5 Can you follow this and try to answer the question?

OpenStudy (anonymous):

ooo so A ???

OpenStudy (mathstudent55):

No. The problem is adding 5x to -3y. Choice A is adding -5x to 3y. They are not the same things. With the commutative property you can change the order of the addition, but you can't change the numbers involved. 5x is not the same as -5x, and 3y is not the same as -3y

OpenStudy (anonymous):

ooo B

OpenStudy (mathstudent55):

right